It shows a 2.0 kg block is being pushed along a rough horizontal surface by a force F=30N at an angle 60° from the normal. If the block moves at constant acceleration 0.5ms-² , what is the coefficient of friction ?
1
Expert's answer
2020-12-14T07:15:51-0500

According to Newton's second law, the interacting forces can be written as


"F\text{ sin}60°-f=ma."

The force of friction here is


"f=F\text{ sin}60°-ma=\mu N,\\\space\\
\mu=\frac{F\text{ sin}60°-ma}{N},"

where the normal force is


"N-mg+F\text{ cos}60°=0,\\
N=mg-F\text{ cos}60°."

So, the coefficient of friction is


"\mu=\frac{F\text{ sin}60°-ma}{mg-F\text{ cos}60°}=5.43."